A bus starting from rest moves with a uniform acceleration of 0.1 m s-2 minutes.
Find
1. the speed acquired
2. the distance traveled.
1
Expert's answer
2011-06-09T08:00:57-0400
Let's use the equations of the motion with constant acceleration.
The speed is:
v=at=0.1*(2*60)=0.1*120=12 m/s
The distance: s=a * t2 / 2 = 0.1* 1202 / 2= 720 meters.